# KAT-V1: Kwai-AutoThink Technical Report

### v2

@@ Abstract
We present Kwaipilot-AutoThink (KAT), an open-source 40B large language model developed to address the overthinking problem in reasoning-intensive tasks, where an automatic thinking training paradigm is proposed to dynamically switch between reasoning and non-reasoning modes based on task complexity. Specifically, first, we construct the dual-regime dataset based on a novel tagging pipeline and a multi-agent synthesis strategy, and then we apply Multi-Token Prediction (MTP)-enhanced knowledge distillation, enabling efficient and fine-grained reasoning transfer with minimal pretraining cost. Besides, we implement a cold-start initialization strategy that introduces mode-selection priors using majority-vote signals and intent-aware prompting. Finally, we propose Step-SRPO, a reinforcement learning algorithm that incorporates intermediate supervision into the GRPO framework, offering structured guidance over both reasoning-mode selection and response accuracy. Extensive experiments across multiple benchmarks demonstrate that KAT consistently matches or even outperforms current state-of-the-art models, including DeepSeek-R1-0528 and Qwen3-235B-A22B, across a wide range of reasoning-intensive tasks while reducing token usage by up to approximately 30%. Beyond academic evaluation, KAT has been successfully deployed in Kwaipilot (i.e., Kuaishou’s internal coding assistant), and improves real-world development workflows with high accuracy, efficiency, and controllable reasoning behaviors. Moreover, we are actively training a 200B Mixture-of-Experts (MoE) with 40B activation parameters, where the early-stage results already demonstrate promising improvements in performance and efficiency, further showing the scalability of the AutoThink paradigm.
@@ 1 Introduction
Through this two-stage paradigm, Kwaipilot-AutoThink not only achieves state-of-the-art performance across multiple benchmarks-including math, code, commonsense, and scientific tasks—but also significantly reduces token consumption. For instance, our model achieves scores of 88.1 and 93.3 on AIME2025 and AIME2024, respectively, while using up to 27.3% fewer tokens than the current state-of-the-art LLMs. Furthermore, Kwaipilot-AutoThink demonstrates robust performance in real-world deployment: it has been integrated into Kwaipilot, Kuaishou’s internal code assistant, where it effectively handles a wide range of development tasks, underscoring the real-world practicality of our thinking control framework and cost-efficient training approach.
@@
  • State-of-the-Art Performance with Efficiency Gains: Kwaipilot-AutoThink achieves state-of-the-art results across multiple benchmarks, spanning math, code, commonsense, and scientific domains. Remarkably, despite having only 40 billion parameters, our model matches or surpasses the performance of leading LLMs with several times more parameters. For example, Kwaipilot-AutoThink scores 88.1 and 93.3 on AIME2025 and AIME2024, respectively, while reducing token usage by up to 27.3% compared to much larger baselines. These results demonstrate that our model delivers both competitive performance and high efficiency, underscoring the effectiveness of our training paradigm, reasoning control framework, and the fine-grained data mixture strategy that balances domain diversity and data complexity.
@@ 4.1.1 Data Construction
To enhance the model’s understanding of reasoning mode selection, we introduce an auxiliary evaluation step. Specifically, for each query-response pair, we use DeepSeek-V3 to generate a judgment rationale explaining the appropriateness of the selected mode. This step provides the training signal necessary for learning the alignment between query complexity and reasoning depth. Futhermore, to ensure coverage and avoid overfitting to biased query distributions, 1% of the queries are randomly assigned a mode (regardless of majority voting), enforcing the model’s exposure to diverse reasoning scenarios.
@@ 4.3 Analysis
We systematically evaluated the evolution of reasoning control (think-on versus think-off activation) and token efficiency across multiple benchmark tasks. Figure 7 shows the proportion of think-on activations across different evaluation datasets during inference. Reasoning-intensive tasks, including AIME2024, AIME2025, GPQA-Diamond, and LiveCodeBench, consistently demonstrate high think-on activation rates (>80%) throughout training. Conversely, tasks that require less explicit reasoning—such as GSM8K, MBPP, HumanEval, CEval, and IFEval—exhibit a clear downward trend in think-on activation as training progresses. Across all datasets, the average think-on rate decreased significantly from approximately 72% to 48%, representing a relative reduction of about 34% and highlighting the model’s enhanced capability for selectively invoking reasoning.
Consistent with reasoning mode trends, Figure 8 illustrates the corresponding dynamics in average token counts per generated response during inference. Datasets with consistently high think-on activation maintain higher average token counts, indicative of sustained explicit reasoning. In contrast, datasets with reduced think-on activation show significant decreases in token usage over the course of training. Specifically, the average number of generated tokens across all datasets decreased from an initial 7,778 to 6,168 tokens , reflecting a relative reduction of approximately 21% and underscoring improvements in inference efficiency achieved through adaptive reasoning control.
@@ 5 Results
In addition to competitive accuracy, KAT demonstrates significant improvements in token efficiency. As shown in Figure 9, across various benchmarks, the average token savings range from 11.6% to 89.9%. Notably, for the 40B model, the average token usage is approximately 72.7% of that used by DeepSeek-R1-0528, showing substantial efficiency gains. These improvements are enabled by KAT’s dynamic reasoning control, which selectively activates reasoning only when necessary, significantly reducing token usage in simpler queries. The model adapts its reasoning depth based on task complexity, resulting in fewer tokens consumed without sacrificing response quality.
@@
MBPP[42] 90.8 93.2 96.2 83.6 65.4 74.6 81.0 72.0
EvalPlus[43, 44] 76.0 80.4 79.2 78.9 77.6 72.1 75.8 67.7
LiveCodeBench[45] 74.2 77.1 73.3 44.7 59.1 61.5 34.1 30.2
BFCL-V3-Mul[46] 39.0 41.6 38.9 29.9 40.1 43.1 16.4 1.9

### v3

@@ Abstract
We present Kwaipilot-AutoThink (KAT), an open-source 40B large language model developed to address the overthinking problem in reasoning-intensive tasks, where an automatic thinking training paradigm is proposed to dynamically switch between reasoning and non-reasoning modes based on task complexity. Specifically, first, we construct the dual-regime dataset based on a novel tagging pipeline and a multi-agent synthesis strategy, and then we apply Multi-Token Prediction (MTP)-enhanced knowledge distillation, enabling efficient and fine-grained reasoning transfer with minimal pretraining cost. Besides, we implement a cold-start initialization strategy that introduces mode-selection priors using majority-vote signals and intent-aware prompting. Finally, we propose Step-SRPO, a reinforcement learning algorithm that incorporates intermediate supervision into the GRPO framework, offering structured guidance over both reasoning-mode selection and response accuracy. Extensive experiments across multiple benchmarks demonstrate that KAT consistently matches or even outperforms current state-of-the-art models, including DeepSeek-R1-0528 and Qwen3-235B-A22B, across a wide range of reasoning-intensive tasks while reducing token usage. Notably, KAT outperforms all open-source models and even surpasses o3-mini on the leakage-controlled LiveCodeBench Pro. Beyond academic evaluation, KAT has been successfully deployed in Kwaipilot (i.e., Kuaishou’s internal coding assistant), where it improves real-world development workflows with high accuracy, efficiency, and controllable reasoning behaviors. Moreover, we are actively training a 200B Mixture-of-Experts (MoE) model with 40B active parameters, and early results already show significant gains, further demonstrating the scalability of the AutoThink paradigm.
@@ 1 Introduction
Through this two-stage paradigm, Kwaipilot-AutoThink not only achieves state-of-the-art performance across multiple benchmarks-including math, code, commonsense, and scientific tasks—but also significantly reduces token consumption. For instance, our model achieves scores of 88.1 and 93.3 on AIME2025 and AIME2024, respectively, while using fewer tokens than the current state-of-the-art LLM. Notably, Kwaipilot-AutoThink ranks first among all open-source models on LiveCodeBench Pro[31], a challenging benchmark explicitly designed to prevent data leakage, and even surpasses strong proprietary systems such as Seed[32] and o3-mini[33]. Furthermore, Kwaipilot-AutoThink demonstrates robust performance in real-world deployment: it has been integrated into Kwaipilot, Kuaishou’s internal code assistant, where it effectively handles a wide range of development tasks, underscoring the real-world practicality of our thinking control framework and cost-efficient training approach.
@@
  • State-of-the-Art Performance with Efficiency Gains: Kwaipilot-AutoThink achieves state-of-the-art results across multiple benchmarks. Remarkably, despite having only 40 billion parameters, our model matches or surpasses the performance of leading LLMs with several times more parameters. For example, Kwaipilot-AutoThink scores leads all open-source models and even outperforms proprietary competitors such as Seed and o3-mini on LiveCodeBench Pro. These results demonstrate that our model delivers both competitive performance and high efficiency, underscoring the effectiveness of our training paradigm, reasoning control framework, and the fine-grained data mixture strategy that balances domain diversity and data complexity.
@@ 4.1.1 Data Construction
To enhance the model’s understanding of reasoning mode selection, we introduce an auxiliary evaluation step. Specifically, for each query-response pair, we use DeepSeek-V3 to generate a judgment rationale explaining the appropriateness of the selected mode. This step provides the training signal necessary for learning the alignment between query complexity and reasoning depth. Furthermore, to ensure coverage and avoid overfitting to biased query distributions, 1% of the queries are randomly assigned a mode (regardless of majority voting), enforcing the model’s exposure to diverse reasoning scenarios.
@@ 4.3 Analysis
We systematically evaluated the evolution of reasoning control (think-on versus think-off activation) and token efficiency across multiple benchmark tasks. Figure 7 shows the proportion of think-on activations across different evaluation datasets during inference. Reasoning-intensive tasks, including AIME2024, AIME2025, GPQA-Diamond, and LiveCodeBench, consistently demonstrate high think-on activation rates (>80%) throughout training. Conversely, tasks that require less explicit reasoning—such as GSM8K, MBPP, HumanEval, CEval, and IFEval—exhibit a clear downward trend in think-on activation as training progresses. Across all datasets, the average think-on rate decreased significantly from approximately 90% to 66%, representing a relative reduction of about 24% and highlighting the model’s enhanced capability for selectively invoking reasoning.
Consistent with reasoning mode trends, Figure 8 illustrates the corresponding dynamics in average token counts per generated response during inference. Datasets with consistently high think-on activation maintain higher average token counts, indicative of sustained explicit reasoning. In contrast, datasets with reduced think-on activation show significant decreases in token usage over the course of training. Specifically, the average number of generated tokens across all datasets decreased from an initial 9,887 to 8,037 tokens , reflecting a relative reduction of approximately 19% and underscoring improvements in inference efficiency achieved through adaptive reasoning control.
@@ 5 Results
In addition to competitive accuracy, KAT demonstrates improvements in token efficiency. As shown in Figure 9, across various benchmarks, token savings are consistently achieved. These improvements are enabled by KAT’s dynamic reasoning control, which selectively activates reasoning only when necessary, significantly reducing token usage in simpler queries. The model adapts its reasoning depth based on task complexity, resulting in fewer tokens consumed without sacrificing response quality.
@@
MBPP[45] 85.2 93.2 96.2 83.6 65.4 74.6 81.0 72.0
LiveCodeBench[46] 73.1 77.1 73.3 44.7 59.1 61.5 34.1 30.2
LCB-Pro-Med[31] 12.7 N/A 7.0 0.0 N/A N/A 0.0 N/A
BFCL-V3-Mul[47] 37.4 41.6 38.9 29.9 40.1 43.1 16.4 1.9